OpenStudy (anonymous):

medal and fan What is the general equation of a circle centered at the point (h, k)

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

is that the equation

OpenStudy (anonymous):

What is the equation of the circle with center at (2, 4) and radius of 6?

OpenStudy (samanthagreer):

The correct answer is ; ( E) (x - 2)² + (y - 4)² = 36.

OpenStudy (samanthagreer):

The correct answer is ; ( E) (x - 2)² + (y - 4)² = 36. Equation of a circle with radius 'r' and center (g, f) is given by, (x - g)² + (y - f)² = r² so, required equation of the circle => (x - 2)² + (y - 4)² = 6² => (x - 2)² + (y - 4)² = 36 OR x² + y² - 4x - 8y - 16 = 0 ---------> (Ans)

OpenStudy (anonymous):

thank you

OpenStudy (samanthagreer):

np can i have a medal please

OpenStudy (anonymous):

sorry i fell asleep

OpenStudy (samanthagreer):

omg lol its ok